OpenStudy (shadyboy):

Select the two square roots of 36 a)6 and -6 b)8 and -8 c)9 and -9 d)12 and -12 d)

OpenStudy (zzr0ck3r):

there is one answer to \(x=\sqrt{36}\) there are two solutions to \(x^2=36\)

OpenStudy (shadyboy):

oh and would it be 9 and -9?

OpenStudy (zzr0ck3r):

what is 6*6=? what is -6*-6 = ?